OpenStudy (anonymous):

Write the equation of the line given the slope and the y-intercept. m = 5, y-intercept (0, -3)

jimthompson5910 (jim_thompson5910):

Use the fact that the equation of any line in slope intercept form is y = mx+b

jimthompson5910 (jim_thompson5910):

So if say the slope was 17 and the y-intercept was (0,8), then y = mx+b y = 17x + b y = 17x + 8 This is NOT the answer as it's just an example

OpenStudy (anonymous):

y=5x+-3 ?

jimthompson5910 (jim_thompson5910):

good, and that simplifies to y = 5x - 3

jimthompson5910 (jim_thompson5910):

I should have pointed out that m is the slope and b is the y-intercept (which is the point (0,b)) But I'm sure you've already figured that part out by now

OpenStudy (anonymous):

So then my answer would be y=5x-3 ?

jimthompson5910 (jim_thompson5910):

yes correct

OpenStudy (anonymous):

Wow, that makes sense. Thanks @jim_thompson5910 !

jimthompson5910 (jim_thompson5910):

glad to be of help

OpenStudy (anonymous):

@jim_thompson5910 I also have a similar problem if you could help....

OpenStudy (anonymous):

Write the equation of the line given the slope and a point on the line. in slope-intercept form, m = -4, point (1, -3)

OpenStudy (anonymous):

Would it be y=-4+-3 ?

jimthompson5910 (jim_thompson5910):

no the trick used above only works if the point looks like (0, ???) where you replace the ??? with any number

jimthompson5910 (jim_thompson5910):

So the trick works if the line goes through (0,6) or (0,19)....but does NOT work for (1,-3)

jimthompson5910 (jim_thompson5910):

you have to do this

jimthompson5910 (jim_thompson5910):

y = mx + b y = -4x + b ... plug in m = -4 -3 = -4(1) + b .... plug in x = 1 and y = -3 (from the point (1,-3)) -3 = -4 + b ... multiply I'll let you finish and solve for b

OpenStudy (anonymous):

-3=-4+-3 ?

jimthompson5910 (jim_thompson5910):

if you have -3 = -4 + b how would you isolate b

OpenStudy (anonymous):

add four to both sides?

OpenStudy (anonymous):

1=b ?

jimthompson5910 (jim_thompson5910):

very good

jimthompson5910 (jim_thompson5910):

b = 1

jimthompson5910 (jim_thompson5910):

so m = -4 and b = 1

jimthompson5910 (jim_thompson5910):

which means y = mx+b turns into y = -4x + 1

OpenStudy (johnweldon1993):

OR you could always use the point slope formula.....(it simplifys to y=mx + b) m = -4, point (1, -3) y - y1 = m(x - x1) you have a point...(1, -3) and a slope (-4) in this equation...m is slope so y - (-3) = -4(x - 1) y + 3 = -4x + 4 finish that simplification :)